# Quantitative Magnetohydrodynamic Modelling of Flux Pumping in ASDEX Upgrade

### v1

@@ 3.4 Nonlinear 3D simulation until quasi-stationary state
The 3D nonlinear simulation is further carried out for a considerably longer timescale of resistive diffusion to achieve the quasi-stationary state of core plasma and observe the identified phenomenon of flux pumping. The time trace of energies in more than 500 ms is plotted in Fig. 8. The system evolves in a rather stationary way, and the dominant mode throughout the simulation is the n = 1 mode.
@@ 4.1 Dynamo term in the toroidal direction and magnetic flux conservation
For the 3D simulation, we split all quantities into an axisymmetric part ()333Note that the subscript 0 on the physical quantities (except , ) in this section denotes the axisymmetric component of the 3D simulation, distinguishing it from the magnetic axis values that appear above. and a non-axisymmetric part (), then the toroidally averaged component of Eq. 5 becomes
@@
The various contributions by different terms on the right-hand side of Eq. 8 are illustrated in Fig. 13 for different times. In all three stages of the 3D simulation, as shown in Fig. 13, the black solid lines show almost the same electric field deficit of 1.6 mV/m. This deficit corresponds to the required electric field that should be provided by an anomalous current diffusion mechanism to maintain the quasi-stationary flat current density profile in the plasma core of the 3D simulation, as shown by Fig. 10 (b). The blue solid lines indicate the negative electric field contributed by the resistivity flattening. The contribution of the resistivity flattening first increases slowly over time and then saturates. This is because as implied by the orange dotted line in Fig. 10 (a), the flux pumping mechanism significantly redistributes the electron temperature in the plasma core and results in a flattened and increased resistivity in the plasma core of the 3D simulation. Besides the resistivity flattening, the dominant mechanism offsetting the driving effect of the current source originates from the nonlinear dynamo term from the non-axisymmetric components, i.e., the 1/1 MHD instability in this study. Specifically, a clear profile of the negative dynamo effect forms in the plasma core and increases over time, as shown by the blue dashed lines in Fig. 13. Nevertheless, we also observed a finite positive contribution on the toroidal electric field in the plasma core from the axisymmetric component of plasma flow, i.e., . It offsets a portion of the negative dynamo from the instability, and eventually results in a discounted negative electric field, which will be referred as the net dynamo, labeled by and red solid lines in Fig. 13.
@@
At the early time point of 90 ms, as shown in Fig. 13 (a), the resistivity flattening effect is ignorable, and the net dynamo amplitude is about 0.3 mV/m in the plasma core, much weaker than the required electric field deficit of 1.6 mV/m. As a result, decreases over time as shown in Fig. 9 (c). At 205 ms as Fig. 13 (b) shows, the resistivity flattening contribution is 0.25 mV/m, while the net dynamo is 1.35 mV/m. Therefore, these two effects play together to balance the electric field deficit generated by the external current drive and result in the saturation of as shown in Fig. 9 (c). At 505 ms of the 3D simulation, the resistivity flattening contribution is 0.3 mV/m, while the net dynamo is 1.6 mV/m. The sum of these two negative contributions is slightly larger than the required electric field deficit, which seems should destroy the quasi-stationary state of the 3D simulation. However, the residues of Eq. 8 [equivalent to ] for different times are plotted in Fig. 14. After 363 ms, we find the residual electric field is almost unaltered until the end of the 3D simulation, exhibiting a remarkably flat profile across a considerable radial range (). The flattened distribution of the residual electric field corresponds to a constant shift of over time in the plasma core, but it will not significantly change the helicity of magnetic field lines and the profile of current density, which are mainly determined by the first or second-order spatial derivatives of . The 3D simulation has been further advanced to 700 ms. It is found that the radial profiles and amplitudes of and remain almost unchanged after 450 ms, indicating the system reaches a balanced stage. On the other hand, not shown here, detailed analyses on the evolution of toroidal magnetic field also demonstrate the conservation of the toroidal magnetic flux at the quasi-stationary stage of the 3D simulation, where a similar cancellation between the two nonlinear coupling terms of [, ] and [, ] is observed in the induction equation of . Note that the resistivity is only applied in the toroidal direction of Eq. 1, thereby restricting the evolutions of , , and within the ideal MHD framework. It can thus be confirmed that the saturation of , as observed in the 3D simulation, is a consequence of the settled radial profile of poloidal magnetic flux and the conservation of the toroidal magnetic flux.
@@ 4.2 The parallel dynamo emf projected to the mean magnetic field
Same as the derivation in Subsec. 4.1, the externally applied boundary loop voltage is neglected, and the physical quantities are divided into their respective axisymmetric () and non-axisymmetric () components. In Eq. 9, the contribution by the axisymmetric plasma flow, i.e., , is naturally excluded since it is perpendicular to the mean magnetic field. Then the toroidally axisymmetric component of Eq. 9 can be rewritten as
where (note that is used for normalization because it has lower numerical noise.). The linear terms in Eq. 10 are eliminated since we are concerned about the axisymmetric components contributing to the dynamo. The nonlinear resistive current diffusion term is negligible as mentioned above. The first term in the last line of Eq. 10 shows the parallel dynamo emf, which is exclusively contributed by the correlated perturbations of plasma flow and magnetic field from the non-axisymmetric MHD instabilities.
@@ 4.3 Current redistribution mechanism from the dynamo effect
The radial distributions of two contributing terms at two different times in the 3D simulation are plotted in Fig. 16. The outer structures () contributed by some tearing instabilities are not the subject of our study, as our primary focus is on the dynamo induced current redistribution within the core region (). The blue solid line indicates the net dynamo term, i.e., . The orange dashed line corresponds to the net current driving effect in the presence of current source and resistive current diffusion, i.e., . The green dash-dotted line plots the sum of these two terms, corresponding to the total current diffusion rate. As shown by Fig. 16 (a), at 90 ms, the positive net current drive is slightly larger than the negative diffusion effect from the net dynamo term in the plasma core, which is consistent with the slow decline of in Fig. 9 (c). Nevertheless, at 505 ms, the excellent cancellation between these two terms is achieved, as shown in Fig. 16 (b), corresponding to the saturation of current density and in the plasma core. In the 2D simulation, the plasma flow is ignorable. Therefore the positive driving effect from the current source term cannot be balanced by the dynamo, resulting in a continuous growth of current density in the plasma core region before the final saturation. Meanwhile, the smoothed current source used in the simulation will result in a less intensive driving effect in the current diffusion equation, since the related term in Eq. 12 is roughly proportional to the second derivative of the current source profile.
The analyses in Fig. 13 and 16 both demonstrate the role of negative net dynamo term in counteracting the driving effect from the current source. The results are consistent with the experimental observation where the sawtooth oscillations are avoided by keeping around unity on the timescale of resistive current diffusion.
@@ 5 Preliminary parameter scan of the current source
In addition to the first current source used in the simulations above [the dashed line in Fig. 17 or Fig. 3 (a)], a preliminary parameter scan of the current source has been conducted to study the influence of its intensity and profile on flux pumping. Specifically, two stronger current sources are designed, as indicated by the dash-dotted and dotted lines in Fig. 17, with maximum intensities ( MA/m2) comparable to the experimental current source. The second current source peaks off-axis at , similar to the slightly off-axis deposition of ECCD in the experiment. The third current source peaks on the magnetic axis to simulate the ideal situation of on-axis ECCD deposition, as more central ECCD deposition results in higher current drive efficiency in experiments (approximately proportional to ) [17]. The initial equilibrium and other parameters remain the same as Sec. 2. However, the resistivity in 2D simulations is increased by two orders of magnitude to reduce the computational time. The 3D simulations are run for over 200 ms until ceases to decrease. The saturated profiles of current density, , and parallel dynamo emf are plotted in Figs. 18 and 19, respectively for the second and third types of current sources.
@@ 6.1 Main conclusions of the present work
In contrast, 3D simulations with realistic resistivity and other relevant parameters first show a fast and under-damped oscillation of due to the fast dynamo effect during the initial saturation stage of the 1/1 quasi-interchange. The fast dynamo effect generates the toroidal electric field with a large amplitude on the order of V/m and modifies the helicity of the mean magnetic field (manifested as the change in q0) on a timescale of a few milliseconds, which is much shorter than that of the resistive diffusion (on the order of seconds). Subsequent nonlinear 3D simulations on the resistive diffusion timescale show that first decreases slightly below unity in the first 200 ms. After that, the system enters into a quasi-stationary stage with saturating around 0.99. The maximum toroidal plasma current density ( 2.5 MA/m2) is maintained close to the equilibrium value with a broad and flat profile for hundreds of milliseconds. The significant difference in the toroidal plasma current density between the 2D and 3D simulations indicates the anomalous current redistribution mechanism in flux pumping, which is equivalent to an additional toroidal electric field deficit of a few mV/m to clamp the current density in the 3D simulation. Detailed analyses show that in the quasi-stationary stage after 200 ms, the slow dynamo effect generated by the 1/1 MHD instability has a comparable but negative amplitude (a few mV/m) in the plasma core, while a limited contribution from the resistivity flattening effect is also observed. The sum of the toroidal electric fields due to the negative dynamo and resistivity flattening is comparable with or even slightly larger than the required electric field deficit. The slight mismatch results in a slow shift of over time but will not obviously change the helicity of the magnetic field and the value of . Because the residue electric field is kept constant in both space and time in the plasma core, and the toroidal magnetic flux conservation is also satisfied in the quasi-stationary stage. On the other hand, the excellent cancellation between the current driving effect from the non-inductive source and the enhanced current diffusion by the slow dynamo effect is confirmed in the 3D simulation, which further validates the effective role of the dynamo in the current redistribution of flux pumping.
@@ 6.2 Outlook for flux pumping modelling
The present MHD simulations still adopt the single fluid, visco-resistive, full MHD equations. The development of the extended MHD model for JOREK will be carried out to introduce other possible dynamo mechanisms, such as the Hall dynamo and diamagnetic dynamo [24, 28, 31], and also to capture more non-ideal corrections to the 1/1 MHD instabilities, such as the two-fluids effect [10, 12], finite Larmor radius (FLR) effect [28], self-consistent bootstrap current evolution [33], energetic particles [47, 48], etc. Ideally, the helical plasma core is expected to self-regulate the amplitudes of various dynamo effects provided by the 1/1 instability in the presence of these corrections to sustain a similar quasi-stationary state with .
@@
All parameter scans of the current source, viscosity, resistivity, etc., and the extended MHD developments are important for understanding the flux pumping mechanism through direct 3D MHD simulations and are also crucial for calibrating a fast surrogate model being developed [49], which aims to efficiently predict the amplitude of dynamo electric field and the feasibility of flux pumping in both existing tokamaks like AUG [17] and JET [21], and future larger devices like ITER and DEMO.
@@
The necessity of the surrogate model [49] arises mainly from the enormous computational demand of initial value simulations. The presented 3D MHD simulation of flux pumping for the AUG discharge involves a significant timescale separation between the resistive diffusion time () and the Alfvén time (). Specifically, the scale separation value of is greater than for AUG. In larger devices, the timescale separation typically increases due to the larger spatial scale, higher electron temperature, and stronger magnetic field. Specifically, with the parameters of the candidate discharges for flux pumping in JET [21], the timescale separation is estimated to be 30% larger than the AUG case. Thus, the 3D MHD simulation for flux pumping in JET is still feasible and is one of the main objectives of future JOREK simulations. However, for typical ITER and DEMO plasmas [14, 4], the timescale separations involved are more than 10 and 40 times larger, respectively, than the AUG situation. Simulations for larger tokamaks also require higher grid resolution. Therefore, 3D nonlinear MHD simulations of flux pumping for ITER and DEMO are impractical based on current computational resources, and the surrogate model [49], after careful calibration, is essential for further assessing flux pumping in these reactor-scale devices.

### v3

@@ 3.4 Nonlinear 3D simulation until quasi-stationary state
The 3D nonlinear simulation is further carried out for a considerably longer timescale of resistive diffusion to achieve the quasi-stationary state of core plasma and observe the identified phenomenon of flux pumping. The time trace of energies over 700 ms is plotted in Fig. 8. The system evolves in a rather stationary way, and the dominant mode throughout the simulation is the n = 1 mode.
@@ 4.1 Dynamo term in the toroidal direction and magnetic flux conservation
For the 3D simulation, we split all quantities into an axisymmetric part ()444Note that the subscript 0 on the physical quantities (except , ) in this section denotes the axisymmetric component of the 3D simulation, distinguishing it from the magnetic axis values that appear above. and a non-axisymmetric part (), then the toroidally averaged component of Eq. 5 becomes
@@
The various contributions by different terms on the right-hand side of Eq. 8 are illustrated in Fig. 13 for different moments. In all three stages of the 3D simulation, as shown in Fig. 13, the black solid lines show almost the same electric field deficit of 1.6 mV/m. This deficit corresponds to the required electric field that should be provided by an anomalous current diffusion mechanism to maintain the quasi-stationary flat current density profile in the plasma core of the 3D simulation, as shown by Fig. 10 (b). The blue solid lines indicate the negative electric field contributed by the resistivity flattening. The contribution of the resistivity flattening first increases slowly over time and then saturates. This is because as implied by the orange dotted line in Fig. 10 (a), the flux pumping mechanism significantly redistributes the electron temperature in the plasma core and results in a flattened and increased resistivity in the plasma core of the 3D simulation. Besides the resistivity flattening, the dominant mechanism offsetting the driving effect of the current source originates from the nonlinear dynamo term from the non-axisymmetric components, i.e., the 1/1 MHD instability in this study. Specifically, a clear profile of the negative dynamo effect forms in the plasma core and increases over time, as shown by the blue dashed lines in Fig. 13. Nevertheless, we also observed a finite positive contribution on the toroidal electric field in the plasma core from the axisymmetric component of plasma flow, i.e., . It offsets a portion of the negative dynamo from the instability, and eventually results in a discounted negative electric field, which will be referred as the net dynamo, labeled by and red solid lines in Fig. 13.
@@
At the early time point of 90 ms, as shown in Fig. 13 (a), the resistivity flattening effect is ignorable, and the net dynamo amplitude is about 0.3 mV/m in the plasma core, much weaker than the required electric field deficit of 1.6 mV/m. As a result, decreases over time as shown in Fig. 9 (c). At 205 ms as Fig. 13 (b) shows, the resistivity flattening contribution is 0.25 mV/m, while the net dynamo is 1.35 mV/m. Therefore, these two effects play together to balance the electric field deficit generated by the external current drive and result in the saturation of as shown in Fig. 9 (c). At 505 ms of the 3D simulation, the resistivity flattening contribution is 0.3 mV/m, while the net dynamo is 1.6 mV/m. The sum of these two negative contributions is slightly larger than the required electric field deficit, which seems should destroy the quasi-stationary state of the 3D simulation. However, the residues of Eq. 8 [equivalent to ] for different moments are plotted in Fig. 14. After 363 ms, we find the residual electric field is almost unaltered until the end of the 3D simulation, exhibiting a remarkably flat profile across a considerable radial range (). The flattened distribution of the residual electric field corresponds to a constant shift of over time in the plasma core, but it will not significantly change the helicity of magnetic field lines and the profile of current density, which are mainly determined by the first or second-order spatial derivatives of . On the other hand, not shown here, detailed analyses on the evolution of toroidal magnetic field also demonstrate the conservation of the toroidal magnetic flux at the quasi-stationary stage of the 3D simulation, where a similar cancellation between the two nonlinear coupling terms of [, ] and [, ] is observed in the induction equation of . Note that the resistivity is only applied in the toroidal direction of Eq. 1, thereby restricting the evolutions of , , and within the ideal MHD framework. It can thus be confirmed that the saturation of , as observed in the 3D simulation, is a consequence of the settled radial profile of poloidal magnetic flux and the conservation of the toroidal magnetic flux.
@@ 4.2 The parallel dynamo emf projected to the mean magnetic field
Same as the derivation in Subsec. 4.1, the externally applied boundary loop voltage and electrostatic potential are neglected, and the physical quantities are divided into their respective axisymmetric () and non-axisymmetric () components. In Eq. 9, the contribution by the axisymmetric plasma flow, i.e., , is naturally excluded since it is perpendicular to the mean magnetic field. Then the toroidally axisymmetric component of Eq. 9 can be rewritten as
where (note that is used for normalization because it has lower numerical noise). The linear terms in Eq. 10 are eliminated since we are concerned about the axisymmetric components contributing to the dynamo. The nonlinear resistive current diffusion term is negligible as mentioned above. The first term in the last line of Eq. 10 shows the parallel dynamo emf, which is exclusively contributed by the correlated perturbations of plasma flow and magnetic field from the non-axisymmetric MHD instabilities.
@@ 4.3 Current redistribution mechanism from the dynamo effect
The radial distributions of two contributing terms at two different moments in the 3D simulation are plotted in Fig. 16. The outer structures () contributed by some tearing instabilities are not the subject of our study, as our primary focus is on the dynamo induced current redistribution within the core region (). The blue solid line indicates the net dynamo term, i.e., . The orange dashed line corresponds to the net current driving effect in the presence of current source and resistive current diffusion, i.e., . The green dash-dotted line plots the sum of these two terms, corresponding to the total current diffusion rate. As shown by Fig. 16 (a), at 90 ms, the positive net current drive is slightly larger than the negative diffusion effect from the net dynamo term in the plasma core, which is consistent with the slow decline of in Fig. 9 (c). Nevertheless, at 505 ms, the excellent cancellation between these two terms is achieved, as shown in Fig. 16 (b), corresponding to the saturation of current density and in the plasma core. In the 2D simulation, the plasma flow is ignorable. Therefore the positive driving effect from the current source term cannot be balanced by the dynamo, resulting in a continuous growth of current density in the plasma core region before the final saturation. Meanwhile, the smoothed current source used in the simulation will result in a less intensive driving effect in the current diffusion equation, since the related term in Eq. 12 is roughly proportional to the second derivative of the current source profile.
The analyses in Figs. 13 and 16 both demonstrate the role of negative net dynamo term in counteracting the driving effect from the current source. The results are consistent with the experimental observation where the sawtooth oscillations are avoided by keeping around unity on the timescale of resistive current diffusion.
@@ 5 Preliminary parameter scan of the current source
In addition to the first current source used in the simulations above [the dashed line in Fig. 17 or Fig. 3 (a)], a preliminary parameter scan of the current source has been conducted to study the influence of its intensity and profile on flux pumping. Specifically, two stronger current sources are designed, as indicated by the dash-dotted and dotted lines in Fig. 17, with maximum intensities ( MA/m2) comparable to the experimental current source. The second current source peaks off-axis at , similar to the slightly off-axis deposition of ECCD in the experiment. The third current source peaks on the magnetic axis to simulate the ideal situation of on-axis ECCD deposition, as more central ECCD deposition results in higher current drive efficiency in experiments (approximately proportional to ) [17]. The initial equilibrium and other parameters remain the same as Sec. 2. However, the resistivity in 2D simulations is increased by two orders of magnitude to reduce the computational time. The 3D simulations are run for over 200 ms until ceases to decrease. The saturated profiles of current density, , and parallel dynamo emf are plotted in Figs. 18 and 19, respectively for the second and third current sources.
@@ 6.1 Main conclusions of the present work
In contrast, 3D simulations with realistic resistivity and other relevant parameters first show a fast and under-damped oscillation of due to the fast dynamo effect during the initial saturation stage of the 1/1 quasi-interchange. The fast dynamo effect generates the toroidal electric field with a large amplitude on the order of V/m and modifies the helicity of the mean magnetic field (manifested as the change in ) on a timescale of a few milliseconds, which is much shorter than that of the resistive diffusion (on the order of seconds). Subsequent nonlinear 3D simulations on the resistive diffusion timescale show that first decreases slightly below unity in the first 200 ms. After that, the system enters into a quasi-stationary stage with saturating around 0.99. The maximum toroidal plasma current density ( 2.5 MA/m2) is maintained close to the equilibrium value with a broad and flat profile for hundreds of milliseconds. The significant difference in the toroidal plasma current density between the 2D and 3D simulations indicates the anomalous current redistribution mechanism in flux pumping, which is equivalent to an additional toroidal electric field deficit of a few mV/m to clamp the current density in the 3D simulation. Detailed analyses show that in the quasi-stationary stage after 200 ms, the slow dynamo effect generated by the 1/1 MHD instability has a comparable but negative amplitude (a few mV/m) in the plasma core, while a limited contribution from the resistivity flattening effect is also observed. The sum of the toroidal electric fields due to the negative dynamo and resistivity flattening is comparable with or even slightly larger than the required electric field deficit. The slight mismatch results in a slow shift of over time but will not obviously change the helicity of the magnetic field and the value of . Because the residue electric field is kept constant in both space and time in the plasma core, and the toroidal magnetic flux conservation is also satisfied in the quasi-stationary stage. On the other hand, the excellent cancellation between the current driving effect from the non-inductive source and the enhanced current diffusion by the slow dynamo effect is confirmed in the 3D simulation, which further validates the effective role of the dynamo in the current redistribution of flux pumping.
@@ 6.2 Outlook for flux pumping modelling
The present MHD simulations still adopt the single fluid, visco-resistive, full MHD equations. The development of the extended MHD model for JOREK will be carried out to introduce other possible dynamo mechanisms, such as the Hall dynamo and diamagnetic dynamo [24, 28, 31], and also to capture more non-ideal corrections to the 1/1 MHD instabilities, such as the two-fluids effect [10, 12], finite Larmor radius (FLR) effect [28], self-consistent bootstrap current evolution [33], etc. Ideally, the helical plasma core is expected to self-regulate the amplitudes of various dynamo effects provided by the 1/1 instability in the presence of these corrections to sustain a similar quasi-stationary state with . In addition to MHD simulations, the MHD-kinetic hybrid model of JOREK [49, 50] will be introduced in future flux pumping research to investigate the synergistic effects between energetic particles and the 1/1 MHD instabilities, including the impact of energetic ions and electrons [51, 52] on the linear stability of the 1/1 MHD instabilities and the amplitude of the nonlinear dynamo term, as well as the spatial and velocity-space redistributions of energetic particles by the MHD instabilities and the dynamo electric field.
@@
The linear stabilities (growth rate and eigenfunction) of the n = 1 MHD mode for the axisymmetric equilibria at the quasi-stationary stage of the 3D JOREK simulations will be further analyzed at different parameters using CASTOR3D (a linear extended visco-resistive MHD code) [53, 54]. When combined with the fine-tuning of the saturated pressure profile and in JOREK nonlinear simulations by parameter scans (current/heating source, viscosity, resistivity, heat conductivity, etc.), this research would allow for the quantitative analyses of different driving effects (pressure gradient and current density) in the dominant MHD mode and the nonlinear dynamo term. All parameter scans and the extended MHD developments are important for understanding the flux pumping mechanism through direct 3D MHD simulations and are also crucial for calibrating a fast surrogate model being developed [55], which aims to efficiently predict the amplitude of dynamo electric field and the feasibility of flux pumping in both existing tokamaks like AUG [17] and JET [21], and future larger devices like ITER and DEMO.
@@
The necessity of the surrogate model [55] arises mainly from the enormous computational demand of initial value simulations. The presented 3D MHD simulation of flux pumping for the AUG discharge involves a significant timescale separation between the resistive diffusion time () and the Alfvén time (). Specifically, the scale separation value of is greater than for AUG, which approximately requires 0.5 million core-hours for a 3D case on Marconi-Fusion supercomputer. In larger devices, the timescale separation typically increases due to the larger spatial scale, higher electron temperature, and stronger magnetic field. Specifically, with the parameters of the candidate discharges for flux pumping in JET [21], the timescale separation is estimated to be 30% larger than the AUG case. Thus, the 3D MHD simulation for flux pumping in JET is still feasible and is one of the main objectives of future JOREK simulations. However, for typical ITER and DEMO plasmas [14, 4], the timescale separations involved are more than 10 and 40 times larger than the AUG situation, respectively. Simulations for larger tokamaks also require higher grid resolution. Therefore, 3D nonlinear MHD simulations of flux pumping for ITER and DEMO are impractical based on current computational resources, and the surrogate model [55], after careful calibration, is essential for further assessing flux pumping in these reactor-scale devices.